OpenStudy (anonymous):

Consider the line -9x-5y=7 . Find the equation of the line that is perpendicular to this line and passes through the point −3, 2 . Find the equation of the line that is parallel to this line and passes through the point −3, 2 .

OpenStudy (mrnood):

it is easier if you convert the original equation into the form y = mx + c can oyu do that first?

OpenStudy (anonymous):

-5y=-9x+7

OpenStudy (mrnood):

no quite - you need to be careful of the - sign

OpenStudy (anonymous):

well can you help me

OpenStudy (mrnood):

I am helping - you are nearly there you just made a simple error when you arranged the formula. Have a look again at what oyu did - be careful of the minus sign....

OpenStudy (anonymous):

5y=9x-7

OpenStudy (mrnood):

no sorry - that is still not correct . slow down a little... -9x-5y=7 first add 9x to each side.... then divide each side by -5

OpenStudy (anonymous):

ok i see so -5y=9x+7 = y= -9/5x+7/5

OpenStudy (mrnood):

ok that's right now this form tells you about the line it is in th e form y=mx +c the value of m is the slope of the line - what is th eslope in this case?

OpenStudy (anonymous):

-9/5

OpenStudy (mrnood):

OHH it's still not right you divide by -5 so it's y= -9/5x -7/5 OK? you are correct about the slope tho - it's -9/5

OpenStudy (mrnood):

so oyu probably know if the slope of a line is m what is the slope of a perpendicular line?

OpenStudy (anonymous):

5/9

OpenStudy (mrnood):

good - that's right so now oyu need to find a line with that slope passing through (-3,2)

OpenStudy (anonymous):

so plug in the coordinates ??

OpenStudy (mrnood):

so oyu know the equation of that line is y = 5/9 x +b and ypou know that when y = 2 then x = -3

OpenStudy (anonymous):

okay now i understand

OpenStudy (mrnood):

yes plug in those values ot solve for b

OpenStudy (anonymous):

2=5/9(-3)+b

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

ugh i -2=5/9(3)+b

OpenStudy (mrnood):

and for the second part the parallel line has the SAME slope as the original , but passes through the point (-3,2) so do the same for that equation

OpenStudy (anonymous):

2=-9/5(-3)+b

OpenStudy (anonymous):

-2 =-9/5(3)+b

OpenStudy (mrnood):

you need to solve for b in both cases

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

how do i do the fraction again

OpenStudy (mrnood):

you wrote this for the first part 2=5/9(-3)+b so 2+(15/9) = b simplify that

OpenStudy (anonymous):

b=11/3

OpenStudy (anonymous):

its asking for the equation of the line so i dont solve it

OpenStudy (mrnood):

correct so the formula is y = mx + b m =5/9, b= 11/3

OpenStudy (anonymous):

y=5/9x+11/3

OpenStudy (mrnood):

that is correct